How do i find the correct answer for II.3 ?


For .2 and l1. 0.2 oz. 3) The weight of extra-large egg has a Normal distribution with a mean of 2 oz and a standard deviation of 1.2. What is the sampling distribution of the mean weight of extra-large egs (i.e, the distribution of the sample mean weight of an egg in a randomly selected carton of a dozen eggs (i.e., 12 eggs))? (3 points) C) N(12, 0.1) m D) N(12, 1) N(2, 0.058) B) N(2, 0.2) Pupulatim) gampe 6-518 In sampling distribution, what is the most feasible probability that an egg in a carton of a dozen eges weighs more than 2.5 oz? (3 points) A) 0.0000 D) 0.2033 B) 0.3119 0.05 2
Answer #1

Solution:-

II.2) (A) The mean weight of an egg in a randomly selected carton of a dozen eggs is N(2, 0.058).

Mean weight of an egg = 2.0

S.D = \frac{\sigma }{\sqrt{n}}

S.D = \frac{0.20}{\sqrt{12}}

S.D = 0.05774

II.3) (A) The most feasible probability that an egg in a carton of a dozen egg weighs more than 2.5 oz is 0.0000.

x = 2.50

By applying normal distribution:-

z = \frac{x-\mu }{\frac{\sigma }{\sqrt{n}}}

z = 8.62

P(z > 8.62) = 0.0000